# Time-IMM: A Dataset and Benchmark for Irregular Multimodal Multivariate Time Series

### v2

@@ Appendix B Limitations
First, the current version of our dataset focuses exclusively on English-language textual data. This choice simplifies preprocessing and alignment but limits the applicability of our benchmark in multilingual or cross-lingual settings, which are common in global-scale systems such as public health surveillance or international finance global_rainfall ; global_trends ; financial_predicting .
Second, our benchmark is currently limited to the task of time series forecasting. We chose this task in part because it does not require additional label collection—making it easier to construct realistic datasets from naturally occurring logs or signals timedrl ; llm4ts_small . Moreover, forecasting serves as a critical upstream task in time series analysis: models that accurately capture future dynamics can support a range of downstream applications including anomaly detection, classification, and decision support run ; ts2vec . Nonetheless, expanding the benchmark to cover these broader tasks would enable more comprehensive evaluation of modeling techniques.
Third, our focus on textual modality reflects the increasing utility of natural language as a rich, unstructured source of auxiliary information. This design decision is motivated by the recent success of large language models (LLMs), which have demonstrated strong capabilities in extracting and contextualizing semantic information from text. However, we do not currently support other modalities such as images, audio, or tabular records—despite their relevance in domains like remote sensing, clinical medicine, or human activity recognition ts_tabular ; ts_audio ; medfuse .
Fourth, we use pretrained LLMs as frozen text encoders and do not explore task-specific fine-tuning. This design simplifies benchmarking and reduces computational cost, but it may underutilize the full potential of the language models, especially in domains where domain-specific adaptation could improve alignment with time series signals calf ; news_to_forecast . To isolate the value of pretrained representations, we conducted additional experiments using Doc2Vec doc2vec as a train-from-scratch alternative. These results, presented in Appendix L.1, confirm that frozen LLMs consistently outperform Doc2Vec, reinforcing their utility even without fine-tuning.
Finally, while we introduce modular mechanisms for timestamp-aware fusion of text and time series, we do not explore more advanced strategies such as retrieval-augmented generation (RAG) timerag ; ts_rag , memory modules optimus ; tsmae , or reasoning ts_reasoning ; chatts , which could further enhance model interpretability and performance.
@@ C.3 Mean Inter-Observation Interval (IOI)
Mean $\text{Mean IOI}=\frac{1}{M}\sum_{i=1}^{M}\frac{\Delta t_{i}}{s}$ (5)
@@ Appendix D Dataset Details and Data Sources
We organize Time-IMM’s nine datasets according to our irregularity taxonomy. For each, we describe the data, point to the source, and summarize how numerical and textual streams were collected.
@@ D.1 Event-Based Logging (Trigger-Based)
GDELT Event Database 2.0 (2021–2024) gdelt_paper . Raw data was downloaded from: http://data.gdeltproject.org/gdeltv2/masterfilelist.txt
@@ D.3 Human-Initiated Observation (Event-Based)
MIMIC-IV clinical database and MIMIC-Notes mimicIV ; MimicNote . Available at: https://physionet.org/content/mimiciv/3.1/
@@ D.4 Operational Window Sampling (Constraint-Based)
FNSPID: Financial News & Stock Prices fnspid .
@@ D.6 Human Scheduling / Availability (Constraint-Based)
StudentLife study (College Experience Dataset) studentlife .
@@
D.7 Missing Data / Gaps (Artifact-Based)
@@
This type of irregularity results from data that is expected but not recorded, often due to reporting failures, technical issues, or data loss. The result is structured gaps in the time series.
@@
ILINet influenza surveillance (CDC) time_mmd .
@@
  • Textual Data: We include weekly CDC public health summaries and provider notes curated from the report and search result collection process introduced by Time-MMD time_mmd . These documents are treated as asynchronous textual observations and provide contextual insight into public health events, seasonal flu trends, and potential reporting disruptions.
@@ D.8 Scheduling Jitter / Delay (Artifact-Based)
CESNET network flow data cesnet .
@@ D.9 Multi-Source Asynchrony (Artifact-Based)
EPA Outdoor Air Quality Sensors epa_air_quality . Numerical time series data was downloaded from: https://aqs.epa.gov/aqsweb/airdata/download_files.html
Textual news articles were retrieved from Common Crawl News, using monthly archives available at: https://data.commoncrawl.org/crawl-data/CC-NEWS/
@@ Appendix E Prompt Designed for LLM Summarization
This appendix provides the prompt templates used for GPT-4.1 Nano to generate textual summaries for each dataset in Time-IMM. All prompts are designed to produce concise, factual, and context-aware outputs within a 5-sentence limit.
@@ E.1 GDELT (Event-Based Logging)
PROMPTS = """
You are an expert in {domain}.
This task is part of building the GDELT 2.0 Event dataset.
Below is a news article:
{article}
Write a concise summary of the information related to {domain}, using no more than 5 sentences.
If the article is not relevant to {domain}, reply with exactly: NA
@@ E.2 RepoHealth (Adaptive or Reactive Sampling)
PROMPTS = """
The following are comments from multiple GitHub issue threads in the repository ’{repo_name}’,
including all historical comments accumulated up to and including those made on a specific day.
{comments}
Write a concise summary of the main technical concerns, questions,
or suggestions raised across these discussions. Focus on key issues, proposed solutions,
and unresolved questions. Your summary should be no more than 5 sentences.
@@ E.3 FNSPID (Operational Window Sampling)
PROMPTS = """
You are an expert in financial markets.
This task is part of building the stock price prediction dataset.
Below is a news article:
{article}
Write a concise summary of the information related to financial markets, using no more than 5 sentences.
If the article is not relevant to financial markets, reply with exactly: NA
@@ E.4 ClusterTrace (Resource-Aware Collection)
PROMPTS = """
You are an expert in large-scale distributed computing and cloud infrastructure.
This task is part of building the Alibaba Cluster Trace Program dataset.
Below is a collection of logs and system behavior summaries:
{article}
Write a concise summary of the system behavior and operational insights relevant to distributed computing, using no more than 5 sentences.
If the content is not relevant to the dataset, or the logs/summaries are not provided, reply with exactly: NA
@@ E.5 EPA-Air (Multi-Source Asynchrony)
PROMPTS = """
You are an expert in weather.
This task is part of building the Air Quality Data Collected at Outdoor Monitors Across the US dataset.
Below is a news article:
{article}
Write a concise summary of the information related to weather, using no more than 5 sentences.
If the article is not relevant to weather, reply with exactly: NA
@@ Appendix H Adapting Regular Time Series Models via Canonical Pre-Alignment
To adapt regular time series forecasting models to irregular data, we follow the canonical pre-alignment strategy introduced in t-PatchGNN t_patchgnn , as illustrated in Figure 7. This method transforms irregularly sampled time series into a fixed-length, regularly spaced representation that can be consumed by standard forecasting architectures.
@@
To help the model reason about timing and missingness, we expand the feature dimension from $N$ to $12N+$ 1. Each input row includes: (i) the observed values or zeros, (ii) the binary mask, and (iii) the normalized timestamp $\tilde{t}_{l}$. This preserves all temporal and structural information in a dense, aligned format.
@@
The forecast queries $Q$ are appended to the end of the input sequence using the same $12N+$ 1 feature schema. Each query timestamp $q_{k}^{(n)}$ is represented by a feature row that includes: (i) zero values for all variable entries (since ground truth is unknown at inference time), (ii) a binary mask set to 0 (indicating no observation—not a forecast target), and (iii) the normalized query timestamp. This mask entry should not be confused with the evaluation-time forecast mask used to compute prediction errors. Including query timestamps in the input allows the model to explicitly condition on the temporal locations of future targets.
@@ I.2 Multimodality Fusion (MMF)
  • Informer informer : 2 encoder layers, 1 decoder layer, attention factor 3.
  • DLinear dlinear : Default configuration without modification.
  • PatchTST patchtst : 1 encoder layer, 1 decoder layer, 2 attention heads.
  • TimesNet timesnet : 2 encoder layers, 1 decoder layer, top-$k=5$, $d_{\text{model}}=16$, $d_{\text{ff}}=32$.
  • TimeMixer timemixer : 2 encoder layers, $d_{\text{model}}=16$, $d_{\text{ff}}=32$, with 3 downsampling layers (window size 2, average pooling).
  • TimeLLM timellm : GPT2 backbone with 6 layers, input/output token lengths 16 and 96, respectively; $d_{\text{model}}=32$, $d_{\text{ff}}=128$.
  • TTM ttm : 3-level attention patching, 3 encoder and 2 decoder layers; $d_{\text{model}}=1024$, decoder dimension 64.
  • CRU cru : Latent state and hidden units set to 32, with square/exponential activations and gravity gates enabled.
  • Latent-ODE latent_ode : Recognition GRU with 32 hidden units, 1 RNN and 1 generator layer, ODE units set to 32.
  • Neural Flow neural_flows : Latent dimension 20, 3-layer hidden network (dimension 32), GRU units 32, coupling flow with 2 layers, time network: TimeLinear.
  • tPatchGNN t_patchgnn : Patch size 24, 1 Transformer layer, 1 GNN layer, 10-dimensional time and node embeddings, $d_{\text{hidden}}=32$.
@@
Appendix L More Experimental Results
This appendix contains two sets of additional results that further validate the effectiveness of multimodal modeling in irregular time series forecasting.
  • Section L.1 investigates the role of the text encoder by replacing large language models (LLMs) with a classical alternative—Doc2Vec doc2vec —and comparing the resulting performance.
  • Section L.2 reports per-dataset comparisons between unimodal (time series only) and multimodal (time series + text) model variants across all nine datasets in Time-IMM.
L.1 Ablation: Replacing LLMs with Doc2Vec
To evaluate the importance of using pretrained language models, we conduct an ablation study replacing frozen LLM-based text encoders with a train-from-scratch alternative—Doc2Vec. Doc2Vec learns fixed-length embeddings for each text input by training a shallow neural model directly on the textual corpus of each dataset, without relying on large-scale pretraining.
Figure 10 summarizes performance across all datasets. We observe that models using frozen LLMs consistently outperform those using Doc2Vec, confirming the benefit of pretrained semantic representations—even in the presence of asynchronous and domain-specific text. These results highlight the importance of strong general-purpose encoders for effective multimodal forecasting under irregular sampling.
L.2 Multimodal vs. Unimodal Performance
To quantify the contribution of textual signals under irregular sampling, we report per-dataset comparisons between unimodal (time series only) and multimodal (time series + text) model variants. Each result reflects the best-performing fusion configuration selected via validation on the corresponding dataset. Performance is measured using mean squared error (MSE), and improvements are reported as relative percentage change from the unimodal baseline.
@@
  • ILINet (Missing Data) — Table 9

### v3

@@ Appendix B Limitations
First, the current version of our dataset focuses exclusively on English-language textual data. This choice simplifies preprocessing and alignment but limits the applicability of our benchmark in multilingual or cross-lingual settings, which are common in global-scale systems such as public health surveillance or international finance [24, 2, 80].
Second, our benchmark is currently limited to the task of time series forecasting. We chose this task in part because it does not require additional label collection—making it easier to construct realistic datasets from naturally occurring logs or signals [9, 11]. Moreover, forecasting serves as a critical upstream task in time series analysis: models that accurately capture future dynamics can support a range of downstream applications including anomaly detection, classification, and decision support [39, 75]. Nonetheless, expanding the benchmark to cover these broader tasks would enable more comprehensive evaluation of modeling techniques.
Third, our focus on textual modality reflects the increasing utility of natural language as a rich, unstructured source of auxiliary information. This design decision is motivated by the recent success of large language models (LLMs), which have demonstrated strong capabilities in extracting and contextualizing semantic information from text. However, we do not currently support other modalities such as images, audio, or tabular records—despite their relevance in domains like remote sensing, clinical medicine, or human activity recognition [33, 34, 23].
Fourth, we use pretrained LLMs as frozen text encoders and do not explore task-specific fine-tuning. This design simplifies benchmarking and reduces computational cost, but it may underutilize the full potential of the language models, especially in domains where domain-specific adaptation could improve alignment with time series signals [42, 66]. To isolate the value of pretrained representations, we conducted additional experiments using Doc2Vec [35] as a train-from-scratch alternative. These results, presented in Appendix M.1, confirm that frozen LLMs consistently outperform Doc2Vec, reinforcing their utility even without fine-tuning.
Finally, while we introduce modular mechanisms for timestamp-aware fusion of text and time series, we do not explore more advanced strategies such as retrieval-augmented generation (RAG) [72, 48], memory modules [37, 22], or reasoning [13, 70], which could further enhance model interpretability and performance.
@@ C.3 Mean Inter-Observation Interval (IOI)
$\text{Mean IOI}=\frac{1}{M}\sum_{i=1}^{M}\frac{\Delta t_{i}}{s}$ (5)
@@ Appendix D Dataset Details and Data Sources
We organize Time‑IMM’s nine datasets according to our irregularity taxonomy. For each, we describe the data, point to the source, and summarize how numerical and textual streams were collected.
@@ D.1 Event-Based Logging (Trigger-Based)
GDELT Event Database 2.0 (2021–2024) [36]. Raw data was downloaded from: http://data.gdeltproject.org/gdeltv2/masterfilelist.txt
@@ D.3 Human-Initiated Observation (Event-Based)
MIMIC‑IV clinical database and MIMIC‑Notes [27, 28]. Available at: https://physionet.org/content/mimiciv/3.1/
@@ D.4 Operational Window Sampling (Constraint-Based)
FNSPID: Financial News & Stock Prices [18].
@@ D.5 Resource-Aware Collection (Constraint-Based)
Alibaba GPU Cluster Trace v2020 [68]. Main repository: https://github.com/alibaba/clusterdata/tree/master?tab=readme-ov-file
@@ D.6 Human Scheduling / Availability (Constraint-Based)
StudentLife study (College Experience Dataset) [46].
@@
D.7 Unplanned Missing Data/Gaps (Artifact-Based)
@@
This irregularity occurs when expected data points are unintentionally missing due to system failures, transmission errors, or reporting issues. Unlike intentional missingness from adaptive sampling or human decisions, these gaps arise without design and are usually treated as random (MCAR or MAR) and handled by imputation.
@@
ILINet influenza surveillance (CDC) [41].
@@
  • Textual Data: We include weekly CDC public health summaries and provider notes curated from the report and search result collection process introduced by Time-MMD [41]. These documents are treated as asynchronous textual observations and provide contextual insight into public health events, seasonal flu trends, and potential reporting disruptions.
@@ D.8 Scheduling Jitter / Delay (Artifact-Based)
CESNET network flow data [32].
@@ D.9 Multi-Source Asynchrony (Artifact-Based)
EPA Outdoor Air Quality Sensors [61]. Numerical time series data was downloaded from: https://aqs.epa.gov/aqsweb/airdata/download_files.html
Textual news articles were retrieved from Common Crawl News, using monthly archives available at: https://data.commoncrawl.org/crawl-data/$C$-NEWS/
@@ Appendix E Prompt Designed for LLM Summarization
This appendix presents the prompt templates used with GPT-4.1 Nano to generate textual summaries for each dataset in Time-IMM. Each dataset employs a single-step, prompt-based pipeline that jointly performs filtering and summarization. During preprocessing, the model generates a concise five-sentence summary only when the input document is topically relevant; otherwise, it returns “NA,” which automatically excludes the document. This unified approach eliminates the need for a separate relevance classifier and ensures consistent semantic filtering across domains. For instance, the prompts remove off-topic news in GDELT, retain technical discussions in RepoHealth, and focus on market-related news in FNSPID. A manual review of 250 sampled summaries showed 100% recall and 88.4% precision in identifying relevant documents.
@@ Appendix H Adapting Regular Time Series Models via Canonical Pre-Alignment
To adapt regular time series forecasting models to irregular data, we follow the canonical pre-alignment strategy introduced in t-PatchGNN [78], as illustrated in Figure 7. This method transforms irregularly sampled time series into a fixed-length, regularly spaced representation that can be consumed by standard forecasting architectures.
@@
To help the model reason about timing and missingness, we expand the feature dimension from $N$ to $2N+1$. Each input row includes: (i) the observed values or zeros, (ii) the binary mask, and (iii) the normalized timestamp $\tilde{t}_{l}$. This preserves all temporal and structural information in a dense, aligned format.
@@
The forecast queries $Q$ are appended to the end of the input sequence using the same $2N+1$ feature schema. Each query timestamp $q_{k}^{(n)}$ is represented by a feature row that includes: (i) zero values for all variable entries (since ground truth is unknown at inference time), (ii) a binary mask set to 0 (indicating no observation—not a forecast target), and (iii) the normalized query timestamp. This mask entry should not be confused with the evaluation-time forecast mask used to compute prediction errors. Including query timestamps in the input allows the model to explicitly condition on the temporal locations of future targets.
@@ I.2 Multimodality Fusion (MMF)
  • Informer [81]: 2 encoder layers, 1 decoder layer, attention factor 3.
  • DLinear [76]: Default configuration without modification.
  • PatchTST [47]: 1 encoder layer, 1 decoder layer, 2 attention heads.
  • TimesNet [69]: 2 encoder layers, 1 decoder layer, top-$k=5$, $d_{\text{model}}=16$, $d_{\text{ff}}=32$.
  • TimeMixer [65]: 2 encoder layers, $d_{\text{model}}=16$, $d_{\text{ff}}=32$, with 3 downsampling layers (window size 2, average pooling).
  • TimeLLM [26]: GPT2 backbone with 6 layers, input/output token lengths 16 and 96, respectively; $d_{\text{model}}=32$, $d_{\text{ff}}=128$.
  • TTM [20]: 3-level attention patching, 3 encoder and 2 decoder layers; $d_{\text{model}}=1024$, decoder dimension 64.
  • CRU [57]: Latent state and hidden units set to 32, with square/exponential activations and gravity gates enabled.
  • Latent-ODE [56]: Recognition GRU with 32 hidden units, 1 RNN and 1 generator layer, ODE units set to 32.
  • Neural Flow [6]: Latent dimension 20, 3-layer hidden network (dimension 32), GRU units 32, coupling flow with 2 layers, time network: TimeLinear.
  • tPatchGNN [78]: Patch size 24, 1 Transformer layer, 1 GNN layer, 10-dimensional time and node embeddings, $d_{\text{hidden}}=32$.
@@
Appendix L Detailed Results on the Effectiveness of Multimodality
To quantify the contribution of textual signals under irregular sampling, we compare unimodal and multimodal model variants across all nine datasets in Time-IMM. Each result reflects the best-performing fusion configuration selected via validation on the corresponding dataset. Performance is measured using mean squared error (MSE), and improvements are reported as relative percentage change from the unimodal baseline.
@@
  • ILINet (Unplanned Missing Data) — Table 9